# with Language Model Diagnostic Assistance

### v1

@@ 6.1 Prediction Accuracy and Time Cost
6.2 Ablation Studies
PLM Backbone. We compare BioGPT to GPT2 at the same size, and a smaller scale model Pythia (Biderman et al., 2023) that has 70 million parameters. BioGPT is based on GPT2 (Radford et al., 2019), but is further pre-trained on a bio-medical corpus. Table 3 shows both model scale and bio-medical pre-training lead to stronger performance.
Linearization Technique. As done in Hegselmann et al. (2023); Manikandan et al. (2023), we replace the true laboratory test name with a standard feature name while keeping the result for that laboratory group unchanged. That is, we linearize laboratory results as follows ( feature1 : value1 |normal-||| feature2 : value2 … ). Table 3 shows using standard feature names leads to better F1 and AUC score as compared to using the raw laboratory test names. One possible reason is that using raw laboratory test names biases ED-Copilot to pick names over informativeness for prediction.
Feature Importance. We examine the change in ED-Copilot performance when training with and without a given feature set from top three popular feature sets: triage information, complete blood count, and chemical group. Table 3 shows removing triage features leads to a significant drop in performance. Our ED collaborators confirmed this made medical sense since initial laboratory tests are based on triage information. Our results also show all laboratory groups are not equally important, with removing the CBC causing a larger drop in performance than excluding the chemical laboratory groups.
Figure 4: Fraction of patients performing lab groups and predicted by ED-Copilot. On average each patient performed 4.7 groups and cost-effective ED-Copilot suggested 2.4 groups.
Hyper-parameters. Hyper-parameters $(\alpha,\beta)$ in Equation 8 control the trade-off between prediction accuracy and time-cost in training. Increasing $\alpha$ trades off sensitivity over specificity, while increasing $\beta$ trades off F1-score over time-cost. Figure 2(a) shows increasing $\alpha$ leads to trade off sensitivity over specificity. Figure 2(b) show increasing $\beta$ causes ED-Copilot to trade off F1-score over time-cost. The hyper-parameters $(\alpha,\beta)$ provide a control for ED clinicians to change ED-Copilot’s behavior to fit their preferences.
6.3 Analysis on Personalized Diagnostic Assistance
Table 4: Critical Outcome prediction accuracy at different cohorts, by the category of lab groups they performed (Top: Performed at least one of the top two most frequent lab groups; Middle: Performed at least one of the middle six lab groups; Rare: Performed one or more rare tests in the last four lab groups). The total number of positive (critical)/negative cases and positive rate is shown in parentheses.
@@ 6.3 Analysis on Personalized Diagnostic Assistance
Personalized treatment strategies are often required since patients have diverse needs and present differently. ED-Copilot learns a personalized representation for each patient and naturally supports personalized diagnostic assistance. On the other hand, previous cost-effective methods based on traditional ML algorithms (e.g., tree-based methods), optimize cost-effectiveness at the full population level, i.e., via feature selection/pruning, rather than personalized for each patient. However, this does not naturally support personalization, unlike ED-Copilot which suggests individualized future tests based on a patient’s triage and previous test results. The following experiments highlight the benefit of personalization via ED-Copilot .
@@
First, we plot both the fraction of patients receiving each group as well as the fraction of patients predicted by ED-Copilot in Figure 4. In the ED, patients undergo a large variety of laboratory groups: after the two most common groups (Chem and CBC), more than half of the patients performed some other tests. ED-Copilot predicts fewer laboratory groups than the observed data since it is trying to minimize time-cost, yet it still predicts a wide variation of recommended tests depending on the patient’s condition. On the other hand, non-personalized cost-effective methods, often withdraw to picking the most frequent groups.
@@
To further understand the benefit of personalization, we partition patients in the test set into three cohorts based on the rarity of the group they were administered; see Table 4 for a precise definition of these cohorts. The middle and rare cohorts have more positive cases, indicating higher severity in these cohorts. We train tree-based baselines using triage information and results from the two most frequent groups (Chem and CBC), and compare it to ED-Copilot. Results in Table 4 show that ED-Copilot and baselines perform comparably on patients in the top cohorts (i.e., patients that present typically). For the middle cohort, ED-Copilot greatly improves F1-score as well. Further, in the rare cohort which has the highest fraction of critical cases as compared to other groups (15% versus 10%), ED-Copilot achieves higher significantly higher sensitivity than other methods. Hence, ED-Copilot provides personalized assistance to provide more equitable care, particularly for high-risk patients.
@@
Figure 5: Comparison of prediction performance on critical outcome of ED-Copilot when restricted or not restricted to tests patients performed in its suggestion at different time cost constraint.
6.4 Unrestricted Lab Group Suggestion
Since MIMIC-ED-Assist is an offline retrospective benchmark, we restrict ED-Copilot during training to only select laboratory groups that patients have received. However, this restriction may not lead to the optimal group suggestion. This experiment investigates ED-Copilot’s performance without this restriction.
Specifically, if ED-Copilot (unrestricted) selects a laboratory group not received by a patient, we impute those results by 00, but add to the time-cost. ED-Copilot (unrestricted) achieves an accuracy of $77\%$ when predicting the next laboratory group, i.e., the group actually received by a given patient, indicating that ED-Copilot (unrestricted) is reasonably good at fitting the clinician’s decision but there are still differences between model suggestions and historical data.
Next, we compare ED-Copilot (unrestricted) to ED-Copilot (restricted) in terms of F1-score for critical outcome when varying the maximum allowed time. Figure 5 show that for low maximum allowed time, ED-Copilot (unrestricted) performs worse than ED-Copilot since if it picks unobserved labs, the cost brings no new information added (no results available). As the number of allowed laboratory groups increases, ED-Copilot (unrestricted) matches the performance of ED-Copilot. Restriction to observed laboratory groups might be necessary for offline evaluation. When used in real scenarios, ED-Copilot would serve as a “Co-Pilot” and suggest laboratory groups from the entire set without restriction, assisting ED clinicians to optimize their workflows.
@@ 7 Conclusion
In this paper, we aim at reducing emergency department wait time through time-cost-effective diagnostic assistance, in collaboration with ED clinicians. Specifically, we curate MIMIC-ED-Assist, a benchmark that contains comprehensive laboratory information annotated with medically relevant groupings alongside key clinical outcomes that serve as useful approximations of ED wait-time. Using MIMIC-ED-Assist, we propose ED-Copilot, an AI system that provides time-cost-effective diagnostic assistance for ED clinicians, by recommending informative laboratory groups, and flagging patients at high risk of critical outcome and lengthened ED stay. ED-Copilot outperforms baselines by significantly improving prediction accuracy and reducing laboratory time costs. Further, its language model backbone allows personalized diagnostic assistance to better address the needs of patients who present atypically. We believe that this work serves as a stepping stone towards AI-driven solutions in the ED, and we hope that MIMIC-ED-Assist will spur interest in applying advancements in ML and language models to tackle this critical healthcare challenge.
8 Acknowledgments
We gratefully acknowledge partial support from NSF grants DMS-2209975, 2015341, NSF grant 2023505 on Collaborative Research: Foundations of Data Science Institute (FODSI), the NSF and the Simons Foundation for the Collaboration on the Theoretical Foundations of Deep Learning through awards DMS-2031883 and 814639, and NIH grant R01GM152718. We also acknowledge support by the Eunice Kennedy Shriver National Institute of Child Health and Human Development of the National Institutes of Health under Award Number K23HD110716. We also thank Yikuan Li from Northwestern University to share their preprocessed sepsis dataset.
Appendix A Triage and Lab Test Group
In Table 5, we include triage assessments and 12 lab groups with 68 tests and their estimated time cost.
@@ Appendix A Triage and Lab Test Group
Inflammation (178 min) Creatine Kinase (CK) C-Reactive Protein
@@ Appendix B Training details
We use a $3$-layer neural network for all MLPs in ED-Copilot. Due to large class imbalance, we use class weights when training the diagnostic predictor $p_{\psi}$. All experiments, training and hyper-parameter tuning are conducted on a NVIDIA RTX A6000 GPU. During the RL phase of training ED-Copilot, we restrict the action space to be on laboratory groups that were administered to patients. That is, we do not select laboratory groups that patients have not received. Performance when this restriction is removed is explored in Section 6.3. To ensure ensure sufficient policy sampling on experience replay buffer and obtain a large batch size, we freeze the weights of our PLM, to train the policy $\pi_{\eta}$, which is parameterized by $\eta$. We use the PPO algorithm (Schulman et al., 2017) to train $\pi_{\eta}$. The loss function employed is as the following for hyper-parameters $c_{1}$, and $c_{2}$:
@@
$\mathcal{L}_{\text{value}}(\eta)$ $=\hat{\mathbb{E}}_{i}\left[(V_{\eta}({\textbf{{h}}}_{<i})-V_{i})^{2}\right],$
Here ${\textbf{{h}}}_{<i}$ is the hidden representation from PLM of last [EOS] in this patient’s linearized sequence in state $s_{i}$. $\mathcal{L}_{\text{clip}}$ is clipped surrogate loss and $\hat{A}_{i}$ is estimated advantages which is regularized by value function $\mathcal{L}_{\text{value}}$. $\text{Entropy}[\pi_{\eta}]$ denotes an entropy term over the states, and $\hat{\mathbb{E}}_{i}$ is the empirical average over the collected dataset. We used a masked actor-critic network with package stable-baseline3 (Raffin et al., 2021). See algorithm 1 for a high-level description of this method.
@@
We list the hyper-parameters in Table 6, including the supervised fine-tuning and reinforcement learning stage. In the RL stage, we use grid-search to tune $\alpha$ and $\beta$ to balance the trade-off between accuracy and cost. The search scope for $\alpha\in\{\frac{1}{8},\frac{1}{4},\frac{1}{2},1,2,4,8,15,16,32,64,256\}$ and $\beta\in\{\frac{1}{100},\frac{1}{50},\frac{1}{20},\frac{1}{10},1,10,100,1000\}$.
The parameters in PPO that are not specified are assigned the default values found in the Python package (Schulman et al., 2017).
@@
The penalty ratio between false positive and lab cost $\beta$ $\frac{1}{100}$
Appendix C Baseline Details
In Table 7, we present tree models’ results with different imputation methods: (1) mean imputation, (2) median imputation, and (3) zero-imputation (i.e., replacing missing values by 0), (4) using a dummy indicator to encode missing values.

### v2

@@ 6.2 Ablation Studies
PLM Backbone. We compare BioGPT to GPT2 which is the same size, and a smaller model Pythia (Biderman et al., 2023) that has 70 million parameters. BioGPT is based on GPT2 (Radford et al., 2019), but is further pre-trained on a bio-medical corpus. Table 3 shows both model scale and bio-medical pre-training lead to stronger performance. We also compare Llama (7B) (Touvron et al., 2023) without RL for this task to BioGPT (345M) without RL, and ED-Copilot. Llama is fine-tuned using Low-Rank Adaptation (LORA) (Hu et al., 2022) for efficiency. Llama 7B (LORA) without RL performs comparably to BioGPT without RL. This indicates that scaling up the backbone parameters is promising avenue to boost absolute performance. However, Llama 7B (LORA) underperforms BioGPT with RL (i.e., ED-Copilot), particularly when it comes to time-cost. This is expected as the PLM backbone and the RL serve different roles in the system. PLM is used to predict outcomes while RL minimizes time-cost. Fine-tuning does not address time-costs. These two aspects are both necessary.
Hyper-parameters. Hyper-parameters $(\alpha,\beta)$ in Equation 8 control the trade-off between prediction accuracy and time-cost in training. Increasing $\alpha$ trades off sensitivity over specificity, while increasing $\beta$ trades off F1-score over time-cost. Figure 3(a) shows increasing $\alpha$ leads to trade off sensitivity over specificity. Figure 3(b) shows increasing $\beta$ causes ED-Copilot to trade off F1-score over time-cost. The hyper-parameters $(\alpha,\beta)$ provide a control for ED clinicians to change ED-Copilot’s behavior to fit their preferences.
Figure 4: Fraction of patients performing laboratory groups and predicted by ED-Copilot. On average each patient performed 4.7 groups and cost-effective ED-Copilot suggested 2.4 groups.
Table 4: Critical Outcome prediction accuracy at different cohorts, by the category of laboratory groups they performed (Top: Performed at least one of the top two most frequent laboratory groups; Middle: Performed at least one of the middle six laboratory groups; Rare: Performed one or more rare tests in the last four laboratory groups). The total number of positive (critical)/negative cases and positive rate is shown in parentheses.
@@
Table 5: Model performance using only triage information under supervised fine-tuning training.
Model Critical Outcome Lengthened ED Stay
F1 AUC Sensitivity Specificity F1 AUC Sensitivity Specificity
Random Forest 0.346 0.767 0.689 0.731 0.181 0.617 0.597 0.571
XGBoost 0.324 0.736 0.673 0.668 0.165 0.603 0.580 0.560
LightGBM 0.341 0.743 0.624 0.738 0.171 0.615 0.554 0.628
3-layer DNN 0.275 0.712 0.611 0.702 0.155 0.553 0.485 0.597
ED-Copilot (SFT) 0.392 0.792 0.708 0.762 0.199 0.675 0.679 0.557
6.3 Analysis on Personalized Diagnostic Assistance
@@
Personalized treatment strategies are often required since patients have diverse needs and present differently. ED-Copilot learns a personalized representation for each patient and naturally supports personalized diagnostic assistance. Previous cost-effective methods based on traditional ML algorithms such as tree-based methods optimize cost-effectiveness at the full population level, e.g., via feature selection. Unlike ED-Copilot, this does not lead to personalized recommendations that are based on triage and previous test results. The following experiments highlight the benefit of personalization via ED-Copilot.
@@
We plot both the fraction of patients receiving each group of tests as well as the fraction of patients predicted by ED-Copilot to undergo each of these groups in Figure 4. This plot shows significant variation in tests received by ED patients. After the two most common groups (CHEM and CBC), more than half of the patients performed some other tests. ED-Copilot reflects this diversity by recommending a variety of tests based on patients’ condition. On the other hand, non-personalized cost-effective methods often withdraw to picking the most frequent groups.
@@
To further understand the benefit of personalized diagnostic assistance, we partition patients in the test set into three cohorts based on the rarity of laboratory groups they were administered. A precise definition of these cohorts can be found in Table 4. The middle and rare cohorts have a larger proportion of positive cases, indicating higher severity. We train tree-based baselines using triage information and results from the two most frequent groups (CHEM and CBC), and compare it to ED-Copilot. Results in Table 6 show ED-Copilot and baselines perform comparably on patients in the low severity cohorts (i.e., patients that present typically). For the middle cohort, ED-Copilot greatly improves F1-score. In the rare cohort which has the highest proportion of critical outcomes (15% versus 10% in middle cohort), ED-Copilot achieves higher significantly higher sensitivity than other methods. This indicates ED-Copilot’s ability to provide personalized assistance which results in more equitable care, particularly for high-risk patients.
Table 6: Sub-group Analysis for Critical Outcome and Lengthened ED Stay.
Model Critical Outcome Lengthened ED Stay
F1 AUC Sensitivity Specificity Avg. Time-cost F1 AUC Sensitivity Specificity Avg. Time-cost
Group: Male, Positive/Negative (Ratio): 155/1329 (10.44%) on critical outcome, 104/1380 (7.01%) on lengthened ED stay
Random Forest 0.387 0.793 0.785 0.692 265 0.207 0.723 0.644 0.704 265
XGBoost 0.379 0.804 0.726 0.732 265 0.258 0.722 0.692 0.659 265
LightGBM 0.391 0.807 0.762 0.742 265 0.223 0.724 0.769 0.597 265
ED-Copilot 0.377 0.780 0.671 0.775 121 0.242 0.670 0.574 0.702 151
Group: Female, Positive/Negative (Ratio): 143/1608 (8.17%) on critical outcome, 127/1624 (7.25%) on lengthened ED stay
Random Forest 0.385 0.819 0.730 0.794 265 0.193 0.677 0.646 0.616 265
XGBoost 0.368 0.826 0.766 0.740 265 0.171 0.645 0.661 0.576 265
LightGBM 0.394 0.813 0.766 0.730 265 0.198 0.683 0.669 0.636 265
ED-Copilot 0.413 0.823 0.727 0.819 129 0.215 0.694 0.713 0.621 160
Group: Age 18-30, Positive/Negative (Ratio): 22/341 (6.06%) on critical outcome, 23/340 (6.34%) on lengthened ED stay
Random Forest 0.277 0.871 0.889 0.838 265 0.197 0.742 0.739 0.738 265
XGBoost 0.413 0.908 0.833 0.884 265 0.306 0.764 0.652 0.824 265
LightGBM 0.390 0.895 0.889 0.913 265 0.259 0.767 0.783 0.629 265
ED-Copilot 0.208 0.849 0.864 0.708 106 0.304 0.823 0.956 0.705 159
Group: Age 31-60, Positive/Negative (Ratio): 97/1169 (7.66%) on critical outcome, 92/1174 (7.27%) on lengthened ED stay
Random Forest 0.367 0.824 0.739 0.794 265 0.196 0.675 0.620 0.670 265
XGBoost 0.367 0.835 0.750 0.775 265 0.228 0.685 0.674 0.658 265
LightGBM 0.400 0.835 0.802 0.756 265 0.213 0.692 0.739 0.611 265
ED-Copilot 0.409 0.800 0.711 0.823 124 0.209 0.697 0.612 0.664 153
Group: Age 61-90, Positive/Negative (Ratio): 189/1261 (13.03%) on critical outcome, 108/1342 (7.45%) on lengthened ED stay
Random Forest 0.411 0.780 0.710 0.753 265 0.208 0.707 0.685 0.621 265
XGBoost 0.383 0.785 0.659 0.789 265 0.188 0.655 0.704 0.552 265
LightGBM 0.405 0.779 0.688 0.757 265 0.192 0.698 0.667 0.622 265
ED-Copilot 0.400 0.749 0.667 0.694 129 0.222 0.695 0.731 0.671 156
6.4 Sub-group Analysis
We compare the performance of ED-Copilot to tree-based baselines across sub-groups including gender and age to study if the algorithm is working fairly across all sub-groups of interest in Table 6. ED-Copilot displays consistent performance, and outperforms baselines across subgroups on both prediction tasks.
6.5 Performance of Unrestricted Lab Group Suggestion
Since MIMIC-ED-Assist is an offline retrospective benchmark, we restrict ED-Copilot during training to only select laboratory groups that patients have received. However, this restriction may lead to sub-optimal group suggestion. We investigate ED-Copilot’s performance without this restriction through the following experiment.
@@ 6.5 Performance of Unrestricted Lab Group Suggestion
Figure 5: Comparison of prediction performance on critical outcome of ED-Copilot when restricted or not restricted to tests patients performed in its suggestion at different time-cost constraints.
If ED-Copilot (unrestricted) selects a laboratory group not received by a patient, we impute those results by 00, but add to the time-cost. ED-Copilot (unrestricted) achieves an accuracy of $77\%$ when predicting the next laboratory group, i.e., the group actually received by a given patient, indicating that ED-Copilot (unrestricted) is reasonably good at replicating clinical decisions but there are still differences between model suggestions and historical data.
Next, we compare ED-Copilot (unrestricted) to ED-Copilot (restricted) in terms of F1-score for critical outcome when varying the maximum allowed time. Figure 5 shows that when maximum allowed time is small, ED-Copilot (unrestricted) performs worse than ED-Copilot. This is because if ED-Copilot (unrestricted) picks a laboratory test that was not administered to a patient, this selection does not result in new information being added. However, as the number of allowed laboratory groups increases, ED-Copilot (unrestricted) matches the performance of ED-Copilot. Restriction to observed laboratory groups is necessary for offline evaluation. In real-world settings, ED-Copilot should serve as a “Co-Pilot” and suggest laboratory groups from the entire set without restriction, and assist ED clinicians in optimizing their workflows.
We also evaluate ED-Copilot on new patients only using triage data, and without any laboratory results which are the information available on patient arrival. We use ED-Copilot only with supervised fine-tuning (SFT), and without RL. Results are displayed in Table 5. Without laboratory results, all models suffer, but ED-Copilot (SFT) outperforms baselines, showing its utility at an early treatment stage.
@@ 7 Conclusion
In this paper, we aim at reducing emergency department wait time through time-cost-effective diagnostic assistance. In collaboration with ED clinicians, we curate MIMIC-ED-Assist, a benchmark that contains comprehensive laboratory information annotated with medically relevant groupings alongside key clinical outcomes that serve as useful approximations to ED wait-time. Using MIMIC-ED-Assist, we propose ED-Copilot, an AI system that provides time-cost-effective diagnostic assistance for ED clinicians, recommending informative laboratory groups and flagging patients at high risks of critical outcome and lengthened ED stay. ED-Copilot outperforms baselines by significantly improved prediction accuracy and reduced laboratory time-costs. Its language model backbone allows for personalized diagnostic assistance to better address the needs of severe patients. We believe this work takes a step towards AI-driven assistance in the ED and hope that MIMIC-ED-Assist spurs interest in applying advancements in AI to tackle this critical healthcare challenge.
Acknowledgments
We gratefully acknowledge partial support from NSF grants DMS-2209975, 2015341, NSF grant 2023505 on Collaborative Research: Foundations of Data Science Institute (FODSI), the NSF and the Simons Foundation for the Collaboration on the Theoretical Foundations of Deep Learning through awards DMS-2031883 and 814639, and NIH grant R01GM152718. We also acknowledge support by the Eunice Kennedy Shriver National Institute of Child Health and Human Development of the National Institutes of Health under Award Number K23HD110716. We also thank Yikuan Li from Northwestern University to share their pre-processed sepsis dataset.
Impact Statement
The development of ED-Copilot marks an advancement in medical technology, offering an AI-driven diagnostic tool to improve patient care in ED. By expediting diagnoses, ED-Copilot has the potential to increase efficiency of ED operations and enhance patient care.
From an ethical perspective, the deployment of ED-Copilot carries a significant responsibility on the privacy and security of patient data. As the system will handle sensitive health information, strict measures, for example, internal-accessible only systems and compliant data protection procedures, must be in place to protect against breaches and misuse, ensuring patient confidentiality.
In the broader societal context, the implementation of ED-Copilot aims to address the issue of ED congestion. It is vital to ensure that the benefits of such technology are accessible to all segments of society, regardless of socioeconomic status. Equity in healthcare technology means that all patients, irrespective of their background, should have the opportunity to benefit from advancements like ED-Copilot. Moreover, the development of ED-Copilot should support healthcare professionals, not replace them, and should be viewed as a tool to assist medical staff, allowing them to focus on the more nuanced aspects of patient care.
Appendix A Triage and Laboratory Test Group
In Table 7, we include triage assessments and 12 laboratory groups with 68 tests and their estimated time-costs.
@@ Appendix A Triage and Laboratory Test Group
Inflammation (INFLAM) (178 min) Creatine Kinase (CK) C-Reactive Protein
@@ Appendix B Training Details
We use a $3$-layer neural network for all MLPs in ED-Copilot. Due to large class imbalance, we use class weights when training the diagnostic predictor $p_{\psi}$. All experiments, training and hyper-parameter tuning are conducted on one NVIDIA RTX A6000 GPU. During the RL phase of training ED-Copilot, we restrict the action space to be on laboratory groups that were administered to patients. That is, we do not select laboratory groups that patients have not received. To ensure sufficient policy sampling on experience replay buffer and obtain a large batch size, we freeze the weights of our PLM. To train the policy $\pi_{\eta}$, we use the PPO algorithm (Schulman et al., 2017). The loss function employed is as the following for hyper-parameters $c_{1}$ and $c_{2}$:
@@
$\mathcal{L}_{\text{value}}(\eta)$ $=\hat{\mathbb{E}}_{i}\left[(V_{\eta}({\textbf{{h}}}_{<i})-V_{i})^{2}\right].$
[…]
@@
The penalty ratio between false positive and laboratory cost $\beta$ $\frac{1}{100}$
We list the hyper-parameters in Table 8, including the supervised fine-tuning and reinforcement learning stage. In the RL stage, we use grid-search to tune $\alpha$ and $\beta$ to balance the trade-off between accuracy and cost. The search scope for $\alpha\in\{\frac{1}{8},\frac{1}{4},\frac{1}{2},1,2,4,8,15,16,32,64,256\}$ and $\beta\in\{\frac{1}{100},\frac{1}{50},\frac{1}{20},\frac{1}{10},1,10,100,1000\}$. The parameters in PPO that are not specified are assigned the default values found in the Python package (Schulman et al., 2017).
Appendix C More Baseline Results
In Table 9, we present tree models’ results with different imputation methods: (1) Mean imputation. (i.e. replacing missing values with the mean value of the non-missing data for the particular feature). (2) Median imputation. (i.e. replacing missing values with the median value of the non-missing data for the particular feature) (3) Zero imputation. (i.e. replacing missing values by 0). (4) A dummy indicator to encode missing values, which is used in XGBoost and LightGBM.